a milk carton has a square base.its height is 11 cm.It has capacity(volume) of 176 cubic centimetres.Find x,the lenght of each side of the base
Answer by edjones(7311)

V=volume, s=side
V=s^2*11=176
s^2*11/11=176/11
s^2=16
s=+-4 sqrt each side
s=4 cm.
.
Ed